# Biomedical SAM 2: Segment Anything in Biomedical Images and Videos

### v1

@@ 4.1.3 Results
As illustrated in Table 1, we conducted a thorough analysis of various SAM 2 variants. The results indicate a clear performance improvement with an increased number of clicks, particularly noticeable in microscopy datasets where the F1 metric improved from 0.0654 to a maximum of 0.3566. Similarly, significant progress was observed in endoscopy datasets, with the DSC score rising from a minimum of 0.4115 to 0.5382, and the NSD score increasing from 0.4227 to 0.5520. Furthermore, we evaluated SAM 2’s performance across two different model sizes. The larger model size outperformed the tiny variant in terms of majority DSC and NSD scores, given the same number of clicks. This suggests that a larger SAM 2 model possesses superior segmentation capabilities.
@@
When comparing the results of zero-shot of SAM 2 to other fine-tuned models specifically designed for medical image segmentation (e.g. nnU-Net), it is evident that the zero-shot performance is inferior. This discrepancy highlights that despite SAM2’s robust transfer learning capabilities, there remains significant room for improvement in the medical imaging domain. This underscores the necessity of fine-tuning of SAM 2 to achieve optimal performance. In addition, it is observed that when SAM 2 employs multiple clicks, SAM 2’s zero-shot results outperform those of SAM, even though SAM utilizes its largest version, SAM_h. It reinforces the advantage of SAM 2’s advanced design and adaptability, compared with SAM.
@@
Table 1 also shows the performance of our proposed method BioSAM 2. Comparative analysis between BioSAM 2 and SAM 2 reveals substantial improvements with BioSAM 2, achieving enhancements ranging from a minimum of 0.0772 to a maximum of 0.5138. This shows BioSAM 2’s superior performance on instruments segmentation and cell segmentation and demonstrates SAM 2’s significant potential in medical imaging domain. SAM 2, being a universal model, requires adaptation, particularly in the medical field, due to its general-purpose design rather than being specialized. Its limited knowledge base on medical datasets and the constrained number of output masks are key factors. While SAM 2 can effectively segment image-level instances, it struggles with accurately segmenting class-level instances. For example, SAM 2 can easily delineate the boundaries of two cells but cannot determine if they belong to the same class. These limitations impact SAM 2’s performance, especially in multi-class medical segmentation datasets. BioSAM 2 thus plays a critical role in bridging these gaps, enhancing the model’s capability to handle the intricacies of medical image segmentation tasks.
@@
Finally, according to Table 1, BioSAM 2 achieved a DSC score of 0.6251 and an NSD score of 0.6427 on the endoscopy dataset. On the microscopy dataset, it obtained an F1 score of 0.5792 and an NSD score of 0.7436. These results surpass most competing methods and are comparable to state-of-the-art models. This performance demonstrates BioSAM 2’s exceptional capability in medical image segmentation, underscoring its potential to deliver high-quality results.
@@ 4.2.1 Datasets
To validate the performance of SAM 2 on biomedical video, we selected two datasets from medical scenarios, which include EndoVis 2018 [47] and EchoNet-Dynamic [48].
@@
Method EndoVis 2018 EchoNet-Dynamic 1-click 3-clicks 5-clicks GT-mask 1-click 3-clicks 5-clicks GT-mask SAM 2 (Hiera-T) 58.4 69.9 71.4 71.6 6.6 56.2 69.8 70.0 SAM 2 (Hiera-S) 59.7 71.3 72.6 73.8 7.3 58.4 70.2 71.5 SAM 2 (Hiera-B+) 60.1 70.8 73.3 74.2 10.0 65.0 71.6 72.8 SAM 2 (Hiera-L) 60.9 71.2 74.8 77.4 7.1 65.9 67.3 73.9
@@
Method EndoVis 2018 EchoNet-Dynamic 1-frame 2-frames 4-frames 6-frames 8-frames 1-frame 2-frames 4-frames 6-frames 8-frames SAM 2 (Hiera-T) 69.9 69.9 72.5 74.9 76.6 56.2 57.0 63.5 68.1 71.2 SAM 2 (Hiera-S) 71.3 71.0 73.0 75.4 77.8 58.4 58.5 64.0 69.5 70.9 SAM 2 (Hiera-B+) 70.8 71.2 74.1 77.6 78.2 65.0 64.9 67.9 70.0 73.4 SAM 2 (Hiera-L) 71.2 73.0 75.2 78.0 79.9 65.9 66.1 70.3 72.0 72.5
@@ 4.2.2 Experimental Setup
In our evaluation of SAM 2 under the condition of zero-shot, we chose Jaccard and F-Score ($\mathcal{J\&F}$) as our evaluation metrics. The Jaccard index describes the Intersection over Union (IoU) between the predicted mask and the ground truth (gt), while the F-Score quantifies the alignment between the boundaries of the predicted mask and the gt boundaries.
@@ 4.2.3 Results
The most significant breakthrough of SAM 2 compared to SAM is its capability to support video tracking of internal objects. As illustrated in Table 2, we conducted a detailed and comprehensive evaluation of the capabilities of SAM 2 in zero-shot segmentation of medical videos. The results indicate that click counts effectively enhance the accuracy of the outcomes. Just a few additional point prompts enable the model to accurately delineate the object boundaries. Notably, the EchoNet Dynamic [48] showed the most significant improvements. As shown in Figure 5, with only one click, SAM 2 segmented the imaging sector area. However, increasing the number of clicks resulted in a substantial rise in the $\mathcal{J\&F}$ scores from single digits to over 70. Moreover, the model generally performs better with larger sizes when the number of clicks remains the same.
Figure 4: The visualization results of SAM 2 in 3-click applications on two medical scenarios.
Figure 5: The failure example of the SAM 2 in 1-click applications on low-resolution ultrasound videos.
A major issue with SAM 2 video segmentation is that if the target object does not appear in the annotated frame, subsequent tracking and segmentation within the video are unfeasible. We addressed this problem by increasing the number of interacted frames. The results in Table 3 demonstrate the enhancements achieved by this method. With the addition of interacted frames, SAM 2 showed improved performance under a 3-click scenario, even surpassing results obtained using only the Ground Truth mask on the first frame. We observed that even if the same object was repeatedly marked in the interacted frames, it still enhanced the subsequent segmentation results. We hypothesize that this injection of information allows the model to better recognize the same object from different perspectives, thereby achieving better outcomes. Overall, while SAM 2 demonstrates some capability in zero-shot segmentation of medical videos, its lack of training in medical content renders it somewhat perplexed in certain medical scenarios. This underscores the importance of training BioSAM 2 on video data in subsequent efforts.
@@ 5 Conclusion
In conclusion, our development of BioSAM 2 represents a significant advancement in biomedical domain. Through the implementation of two specialized evaluation pipelines designed for single-frame biomedical images and multi-frame biomedical videos, we have rigorously assessed the performance of SAM 2 across diverse medical scenarios and objects of interest. Our results indicate that the enhanced BioSAM 2 not only outperforms current state-of-the-art foundation methods, but also surpasses the performance of majority specialized models trained specifically for the same medical modalities. These findings affirm the potential of BioSAM 2 as a novel biomedical segmentation approach towards more efficient, accurate, and adaptable diagnostic technologies.

### v2

@@ 4.1.3 Results
As illustrated in Table 1 and Table 2, we conducted a thorough analysis of various SAM-2 variants. The results indicate a clear performance improvement with an increased number of clicks, particularly noticeable in microscopy datasets where the F1 metric improved from 0.0654 to a maximum of 0.3566. Similarly, significant progress was observed in other datasets. For instance, the DSC score of endoscopy dataset rises from a minimum of 0.4115 to 0.5382, and the NSD score increases from 0.4227 to 0.5520. Furthermore, we evaluated SAM-2’s performance across two different model sizes. The larger model size outperformed the tiny variant in terms of majority DSC and NSD scores, given the same prompt. This suggests that a larger SAM-2 model possesses superior segmentation capabilities.
@@
When comparing the results of zero-shot of SAM-2 to other fine-tuned models specifically designed for medical image segmentation (e.g. nnU-Net), it is evident that the zero-shot performance is inferior. This discrepancy highlights that despite SAM2’s robust transfer learning capabilities, there remains significant room for improvement in the medical imaging domain. This underscores the necessity of fine-tuning of SAM-2 to achieve optimal performance. In addition, it is observed that when SAM-2 employs proper prompts, SAM-2’s zero-shot results outperform those of SAM, even though SAM utilizes its largest version, SAM_h. It reinforces the advantage of SAM-2’s advanced design and adaptability, compared with SAM.
@@
Table 1 and Table 2 also show the performance of our proposed method BioSAM-2. Comparative analysis between BioSAM-2 and SAM-2 reveals substantial improvements with BioSAM-2, achieving enhancements ranging from a minimum of 0.0772 to a maximum of 0.5138. This shows BioSAM-2’s superior performance on biomedical imaging segmentation and demonstrates SAM-2’s significant potential in this domain. SAM-2, being a universal model, requires adaptation, particularly in the medical field, due to its general-purpose design rather than being specialized. Its limited knowledge base on medical datasets and the constrained number of output masks are key factors. While SAM-2 can effectively segment image-level instances, it struggles with accurately segmenting class-level instances. For example, SAM-2 can easily delineate the boundaries of two cells but cannot determine if they belong to the same class. These limitations impact SAM-2’s performance, especially in multi-class medical segmentation datasets. BioSAM-2 thus plays a critical role in bridging these gaps, enhancing the model’s capability to handle the intricacies of medical image segmentation tasks.
@@
Finally, according to Table 1, BioSAM-2 achieved a DSC score of 0.6251 and an NSD score of 0.6427 on the endoscopy dataset. On the microscopy dataset, it obtained an F1 score of 0.5792 and an NSD score of 0.7436. These results surpass most competing methods and are comparable to state-of-the-art models. This performance demonstrates BioSAM-2’s exceptional capability in medical image segmentation, underscoring its potential to deliver high-quality results.
Table 2: Results summary of 3D organ segmentation on abdomen CT and MR datasets.
Methods Organs in Abdomen CT Organs in Abdomen MR DSC NSD DSC NSD nnU-Net 0.8615$\pm$0.0790 0.8972$\pm$0.0824 0.8309$\pm$0.0769 0.8996$\pm$0.0729 SegResNet 0.7927$\pm$0.1162 0.8257$\pm$0.1194 0.8146$\pm$0.0959 0.8841$\pm$0.0917 UNETR 0.6824$\pm$0.1506 0.7004$\pm$0.1577 0.6867$\pm$0.1488 0.7440$\pm$0.1627 SwinUNETR 0.7594$\pm$0.1095 0.7663$\pm$0.1190 0.7565$\pm$0.1394 0.8218$\pm$0.1409 U-Mamba 0.8638$\pm$0.0908 0.8980$\pm$0.0921 0.8501$\pm$0.0732 0.9171$\pm$0.0689 SAM 0.2957$\pm$0.2761 0.2967$\pm$0.3018 0.3710$\pm$0.4321 0.3712$\pm$0.4321 SAM-2 (Hiera-T) 0.4431$\pm$0.2690 0.3875$\pm$0.2544 0.5534$\pm$0.2609 0.5572$\pm$0.2683 SAM-2 (Hiera-L) 0.4744$\pm$0.2472 0.4208$\pm$0.2306 0.5492$\pm$0.2315 0.5526$\pm$0.2341 BioSAM-2 (Hiera-T) 0.7632$\pm$0.0836 0.8568$\pm$0.0732 0.7439$\pm$0.0629 0.8317$\pm$0.0746
@@ 4.2.1 Datasets
To validate the performance of SAM-2 on biomedical video, we selected two datasets from medical scenarios, which include EndoVis 2018 [51] and EchoNet-Dynamic [52].
@@
Method EndoVis 2018 EchoNet-Dynamic 1-click 3-clicks 5-clicks GT-mask 1-click 3-clicks 5-clicks GT-mask SAM-2 (Hiera-T) 58.4 69.9 71.4 71.6 6.6 56.2 69.8 70.0 SAM-2 (Hiera-S) 59.7 71.3 72.6 73.8 7.3 58.4 70.2 71.5 SAM-2 (Hiera-B+) 60.1 70.8 73.3 74.2 10.0 65.0 71.6 72.8 SAM-2 (Hiera-L) 60.9 71.2 74.8 77.4 7.1 65.9 67.3 73.9
@@
Method EndoVis 2018 EchoNet-Dynamic 1-frame 2-frames 4-frames 6-frames 8-frames 1-frame 2-frames 4-frames 6-frames 8-frames SAM-2 (Hiera-T) 69.9 69.9 72.5 74.9 76.6 56.2 57.0 63.5 68.1 71.2 SAM-2 (Hiera-S) 71.3 71.0 73.0 75.4 77.8 58.4 58.5 64.0 69.5 70.9 SAM-2 (Hiera-B+) 70.8 71.2 74.1 77.6 78.2 65.0 64.9 67.9 70.0 73.4 SAM-2 (Hiera-L) 71.2 73.0 75.2 78.0 79.9 65.9 66.1 70.3 72.0 72.5
@@ 4.2.2 Experimental Setup
In our evaluation of SAM-2 under the condition of zero-shot, we chose Jaccard and F-Score ($\mathcal{J\&F}$) as our evaluation metrics. The Jaccard index describes the Intersection over Union (IoU) between the predicted mask and the ground truth (gt), while the F-Score quantifies the alignment between the boundaries of the predicted mask and the gt boundaries.
@@ 4.2.3 Results
The most significant breakthrough of SAM-2 compared to SAM is its capability to support video tracking of internal objects. As illustrated in Table 3, we conducted a detailed and comprehensive evaluation of the capabilities of SAM-2 in zero-shot segmentation of medical videos. The results indicate that click counts effectively enhance the accuracy of the outcomes. Just a few additional point prompts enable the model to accurately delineate the object boundaries. Notably, the EchoNet Dynamic [52] showed the most significant improvements. As shown in Figure 5, with only one click, SAM-2 segmented the imaging sector area. However, increasing the number of clicks resulted in a substantial rise in the $\mathcal{J\&F}$ scores from single digits to over 70. Moreover, the model generally performs better with larger sizes when the number of clicks remains the same.
Figure 4: The visualization results of SAM-2 in 3-click applications on two medical scenarios.
Figure 5: The failure example of the SAM-2 in 1-click applications on low-resolution ultrasound videos.
A major issue with SAM-2 video segmentation is that if the target object does not appear in the annotated frame, subsequent tracking and segmentation within the video are unfeasible. We addressed this problem by increasing the number of interacted frames. The results in Table 4 demonstrate the enhancements achieved by this method. With the addition of interacted frames, SAM-2 showed improved performance under a 3-click scenario, even surpassing results obtained using only the Ground Truth mask on the first frame. We observed that even if the same object was repeatedly marked in the interacted frames, it still enhanced the subsequent segmentation results. We hypothesize that this injection of information allows the model to better recognize the same object from different perspectives, thereby achieving better outcomes. Overall, while SAM-2 demonstrates some capability in zero-shot segmentation of medical videos, its lack of training in medical content renders it somewhat perplexed in certain medical scenarios. This underscores the importance of training BioSAM-2 on video data in subsequent efforts.
@@ 5 Conclusion
In conclusion, our development of BioSAM-2 represents a significant advancement in biomedical domain. Through the implementation of three specialized evaluation pipelines designed for biomedical images and videos, we have rigorously assessed the performance of SAM-2 across diverse medical scenarios and objects of interest. Our results indicate that the enhanced BioSAM-2 not only outperforms current state-of-the-art foundation methods, but also surpasses the performance of majority specialized models trained specifically for the same medical modalities. These findings affirm the potential of BioSAM-2 as a novel biomedical segmentation approach towards more efficient, accurate, and adaptable diagnostic technologies.